OpenStudy (anonymous):

Find the coordinates of the other endpoint of a segment with the given endpoint and mid point 28) Q(-5, 9) M(-8, -2) 29) A(6,7) M(10, -7) 30) D(-3.5, -6) M(1.5, 4.5)

OpenStudy (anonymous):

I'd just extrapolate it out. Or did they give you formulas they wanted you to use? 28) Change from start to midpoint: -8-(-5) = -3 -2-9 = -11 28) End point -8-3 = -11 -2-11 = -13 Therefore, the unknown endpoint is (-11,-13) for Q28

OpenStudy (anonymous):

29) (14,-21) 30) (6.5,15)